Please help me out with this I have NO idea how to begin solving this problem.
The sum of two numbers is 100. One number is at least 100. One number is at least 16 more than the other number. What are the two numbers?

Answer by Paul(988)

Let smaller number be x
Let the larger number be x+16
(x)+(x+16)=100
2x+16=100
2x=84
x=42
42+16=58
Hence, one number is 42 and another one is 58.
Paul.